OpenStudy (anonymous):

x^2-2x+3=0.. how many real solutions?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

infinitely many solutions

OpenStudy (anonymous):

False

OpenStudy (anonymous):

?

OpenStudy (anonymous):

That's not true

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

then one solution if my answer isn't true

OpenStudy (anonymous):

but my answer is true

jimthompson5910 (jim_thompson5910):

Use the discriminant D = b^2 - 4ac

OpenStudy (anonymous):

it's infinitely many solutions simple algebra bro

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{2a}\] ax^2+bx+c a=1, b=-2, c=3

OpenStudy (anonymous):

there r infinitely many solutions

OpenStudy (anonymous):

That's not even a possibility

OpenStudy (anonymous):

Is it 2 @AJ01?

OpenStudy (anonymous):

sorry i was a way...\[\frac{ -(-2)\pm \sqrt{(-2)^{2}-4*1*2} }{ 1*2 }\]

OpenStudy (anonymous):

\[\frac{ 2\pm \sqrt{4-12} }{ 2 }=1\pm \frac{ \sqrt{-8} }{ 2}=1\pm \frac{ \sqrt{4*-2} }{ 2}\]

OpenStudy (anonymous):

\[1\pm \sqrt{-2}\]

jimthompson5910 (jim_thompson5910):

AJ01, you plugged in c = 2 when it should be c = 3

OpenStudy (anonymous):

thanks but the multiplication is 3*4=12 just miss typing